# Matrix Multiplication in the MPC Model

### v1

@@ Abstract
In this paper, we study the matrix multiplication problem in the MPC model. We have two matrices, and the task is to compute their product. These matrices are evenly distributed over $P$ processors. Each processor has $M$ memory such that $P\cdot M\geq$ (size of the matrices). The computation proceeds in synchronous rounds. In a communication round, a processor can send and receive messages to(from) any other processor, with the total size of messages sent or received being $O(M)$. We give an almost complete characterisation of the problem in various settings. We prove tight upper bounds and lower bounds for the problems in three different settings–when the given input matrices are (i) general square matrices, (ii) rectangular matrices, and (iii) sparse square matrices (that is, each row and column contains a bounded number of nonzero elements). In particular, we prove the following results:
@@
    Multiplication of two $n\times n$ matrices in the MPC model with $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, requires $\Theta(n^{\frac{\alpha}{2}})$ rounds in semirings.
@@
    Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ (where $d\leq n$) respectively, with $n$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings.
@@
    Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ ( where $d\leq n$) respectively requires
    1. (a)
      $\Theta(\sqrt{d}+\log_{d}n)$ rounds with $n$ processors and $O(d)$ memory per processor in semirings
    2. (b)
      $\Theta(\frac{d}{\sqrt{n}})$ rounds with $d$ processors and $O(n)$ memory per processor in semirings.
  4. 4.
    Multiplication of two $d$-sparse matrices (each row and column of the matrices contains at most $d$-nonzero elements) with $n$ processors and $O(d)$ memory per processor can be done in $O(d^{0.9})$ rounds in semirings.
keywords:
Massively parallel computing, distributed algorithms, , matrix multiplication
category:
[…]
@@ 1 Introduction
We are given two matrices $A$ and $B$, and our task is to compute $C=A\times B$. The complexity to solve this problem differs with the dimensions and structure of the input matrices. Therefore, we explore the problem for the following cases that are often studied: when matrices $A$ and $B$ are (i) square matrices i.e. both of dimension $n\times n$, (ii) rectangular matrices of dimension $n\times d$ and $d\times n$, (iii) rectangular matrices of dimension $d\times n$ and $n\times d$, (iv) $d$- sparse matrices i.e. both are $n\times n$ size matrices such that each matrix contains most $d$ non-zero elements in each row and each column. The problem has been studied for all these cases in various models of distributed and parallel algorithms, for example, the clique model, the low bandwidth model etc. [6, 3, 2, 1, 4, 5]. We study this problem in the MPC model.
@@ 1.1 MPC Model
In models like clique and the low bandwidth, the primary restriction is put on the amount of communication between the nodes, whereas the internal computation within a node (or processor) remains unbounded in terms of time and space. In the MPC model, local memory is the primary bottleneck. Consider an input problem of size $N$. To solve the problem in the MPC model, we have $P$ processors, each with $M$ memory. The input is distributed among $P$ processors. Therefore, we have the following natural constraints: $P\cdot M\geq N$ and $P\cdot M\geq Output-size$ in order to store the complete input and output, respectively. The computation proceeds in synchronous rounds. Processors perform computation locally, and at the end of a round, they exchange their messages. Because of the memory restriction, each processor is allowed to send and receive $O(M)$ amount of information in each round. If any processor receives more than $O(M)$ amount of information during the course of any algorithm in the MPC model, the algorithm is considered failed.
@@ 1.2 Setup
In this paper, we primarily focus on the matrix multiplication problem in semirings. We assume that initially, matrices $A$ and $B$ are distributed evenly among all the processors. The structure of the distribution does not matter because we can reach any desired even distribution in one round (even if we start with an uneven distribution). At the end of the algorithm the elements of the output matrix $C$ should also be evenly distributed. We analyse the problem in the following three settings.
  1. 1.
    When $A$ and $B$ are square matrices of size $n\times n$. We call this $(n,n,n)$ case.
  2. 2.
    $A$ and $B$ are rectangular matrices. We divide it into the following two cases
    1. (a)
@@
    2. (b)
      $(d,n,d):A$ and $B$ are of size $d\times n$ and $n\times d$ respectively
  3. 3.
    $A$ and $B$ are $d$-sparse $n\times n$ matrices, i.e. each row and column of matrices $A$ and $B$ contains at most $d$ non-zero elements. We call it the $d$-sparse case.
In this paper, we also assume that $P.M\leq O(N)$ i.e. the total memory of the system is just tight enough to hold the input and output matrices.
@@ 1.3 Our Results
We prove that $(n,d,n)$ case can be solved in $O(n^{\frac{\alpha}{2}})$ rounds. Our algorithm that gives this bound is very straightforward. However, what is interesting is that we prove that this simple algorithm is the best we can hope for, i.e., we prove the matching lower bound. We prove that it will take $\Omega(n^{\frac{\alpha}{2}})$ rounds to solve the problem. For $(n,d,n)$ case, we prove that the problem can be solved in $\Theta(\frac{d}{\sqrt{n}})$ rounds with $n$ processors each with $O(n)$ memory. Similarly, for the $(d,n,d)$ case, we prove that it will take $\Theta(\sqrt{d}+\log_{d}n)$ round when number of processors are $n$, and $\Theta(\frac{d}{\sqrt{n}})$ rounds when number of processors are $d$. For the $d$-sparse case, when the number of processors is $n$, we improve the trivial bound of $O(d)$ to $O(d^{0.9})$. In this case, we could not prove a lower bound, though we notice that some instances of this type can be solved in $O(\sqrt{d})$ rounds (when all the elements are concentrated along the diagonal in the input matrices).
1.3.1 High-level Idea of Our Technique
For square and rectangular cases, we observed that in order to reduce inter-processor communication, the best case is when all processors try to compute a square submatrix of $C$. Thus, we divide the matrices $A$ and $B$ into smaller square matrices of equal size depending on the local memory of the processor, and assign each submatrix to one of the processors. At the end, each processor needs to produce the corresponding submatrix of $C$. The interesting part lies in the lower bound. Since we are working with semirings, we know that we need to compute a total $n^{3}$ products (in case of $(n,n,n)$, and $n^{2}d$ and $d^{2}n$ products in case of $(n,d,n)$ and $(d,n,d)$ respectively) of type $a_{ik}b_{kj}$, where $a_{ik}$ and $b_{kj}$ are the elements of matrices $A$ and $B$ respectively. We analyse the maximum number of products (we call it a term in this paper) a processor can compute in each round. Since we are working with semirings, in order to compute an element of matrix $C$, a processor needs to have information about all the terms required to compute that element explicitly. We show that in $(n,n,n)$ case, because of the memory restriction of $O(n^{2-\alpha})$, a processor cannot compute more than $O(n^{\frac{3}{2}\alpha})$ terms in one round. Since there are $n^{\alpha}$ processors, it will take $\Omega(n^{\frac{\alpha}{2}})$ rounds. A similar argument works for rectangular cases.
@@ 1.3.1 High-level Idea of Our Technique
For the $d$-sparse case, we extend the idea of [4]. In their paper, they show that $d$-sparse matrix multiplication is equivalent to processing triangles in a $d$-degree graph. In other words, processing triangles is equivalent to computing terms. They show that given any $d$-degree graph, say $T$, we can divide the graph into two parts such that one part, say $T^{\prime}$, is clustered - it contains layers of subgraphs of the original graph such that in each layer we have disjoint clusters of triangles which are small and dense, and the other part, say $T^{\prime}$, contains a low number of triangles. Now, in order to process all the triangles, we can first process all the triangles in the clustered part layer-by-layer using square matrix multiplication, then compute all the triangles $T^{\prime}$ using brute force. This is a graph-theoretical result independent of any model. In this paper, we refined their result (we will describe it in a later section) and then use our square matrix multiplication as a black box to solve the clustered part.
2 Square Matrix Multiplication
In this section, we will prove that if there are $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, then $(n,n,n)$ case can be solved in $O(n^{\alpha/2})$ rounds in the MPC model. This required memory is tight, as the size of input and output is $O(n^{2})$. For the sake of simplicity, we first give the analysis of the case when $\alpha=1$. Then we prove the result in full generality.
Theorem 2.1.
Matrix multiplication in MPC model with $n$ processors, each with $O(n)$ memory, can be performed in $O(\sqrt{n})$ rounds for semirings.
Proof 2.2.
We shall prove this by proposing a simple algorithm which achieves $O(\sqrt{n})$ bound. Divide each of the matrices $A$, $B$ and $C$, into $n$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$. Let $M^{ij}$ denote the sub-matrix which consists of elements of $M$ at rows $(i-1)\sqrt{n}+1,(i-1)\sqrt{n}+2,\ldots,i\sqrt{n}$ and columns $(j-1)\sqrt{n}+1,(j-1)\sqrt{n}+2,\ldots,j\sqrt{n}$ (where $M\in\{A,B,C\})$ Let us index the $n$ processors as $p_{11},p_{12},\ldots,p_{\sqrt{n}\sqrt{n}}$. Assign the submatrices $A^{ij}$, $B^{ij}$ and $C^{ij}$ to processor $p_{ij}$. Initialize $C^{ij}$ as a null matrix at $p_{ij},\forall\ i,\ j$. The algorithm works as follows.
for $x\leftarrow 1$ to $\sqrt{n}$ do
       send all elements of $A^{ix}$ and $B^{xj}$ to $p^{ij}$ $\forall\ i,\ j\leq\sqrt{n}$
$p_{ij}$ locally computes $C^{ij}$ = $C^{ij}$ + $A^{ix}*B^{xj}$ $\forall\ i,\ j$ and * denotes matrix multiplication
(Note that this step can be done without exceeding $O(n)$ memory)
Algorithm 1 Square Matrix Multiplication in $O(\sqrt{n})$ rounds
The above algorithm follows all the constraints of the MPC model. It is easy to see that during the course of the algorithm, each node sends at most $n\sqrt{n}$ elements to other nodes and receives at most $n\sqrt{n}$ elements from other nodes. Since each element can send and receive $O(n)$ elements in each round, the above algorithm will terminate in $O(\sqrt{n})$ rounds. At the end, processor $p_{ij}$ will contain the submatrix $C^{ij}$ of the final output matrix $C$. This finishes the proof of Theorem 2.1.
Now we will proceed towards proving the lower bound and prove that the simple algorithm presented above is actually tight and cannot be improved further. Before that, we will prove the following lemma, which we will use to prove all the lower bounds that we present in this paper.
Lemma 2.3.
A processor with $r$ memory can calculate at most $O(r\sqrt{r})$ terms in one round in the MPC model for any matrix multiplication instance.
Proof 2.4.
We prove the lower bound by finding the limit on computational capacity per processor. We know that $c_{ij}$ = $\sum_{x=1}^{n}a_{ix}*b_{xj},$ $\forall\ i,\ j$. As mentioned earlier, we refer to $a_{ix}*b_{xj}$ as one term of matrix $C$ (since it is used to compute an element of $C$). We try to find the maximum number of terms that a processor can calculate in one round. Let us assume that $t$ is the maximum number of terms of $C$ that a processor can calculate in one round. Each of these terms is of the form $a_{ix}*b_{xj}$. We segregate the terms into groups on the basis of the values of $x$, so that the terms with the same value of $x$ are in the same group. Let $H_{1},H_{2},\ldots$ be different groups containing $t_{1},t_{2},t_{3},...$ terms respectively (see Figure 1). We try to find the minimum memory required by a processor to calculate these terms. We know that in the case of semirings, in order to compute a term, say $a_{ix}*b_{xj}$, a processor needs to receive both elements $a_{ix}$ and $b_{xj}$ distinctly. Notice that no elements of $A$ and $B$ are common between the terms present in any two different groups. So we try to find the minimum memory required to store the elements of $A$ and $B$ involved in the terms of group $H_{i}$, and sum it for all values of $i$. This sum should not exceed the $O(n)$ memory available with a processor. Let $p$ and $q$ be the number of elements of $A$ and $B$ required to calculate the terms of $H_{i}$. We can say that, $p*q\geq t_{i}$ (since each term is unique, $t_{i}$ combinations should exist). This implies that $p+q\geq 2\sqrt{t_{i}}$. Thus, at least $2\sqrt{t_{i}}$ memory is required per group, leading to the total memory requirement of $2\sum\sqrt{t_{i}}\leq O(r)$. Notice that the processor computing terms of $H_{k}$ needs to keep the value of all the terms (i.e the result of the product $a_{ik}b_{kj})$ present in $H_{i}$ in different memory locations. Because any two terms, say $a_{i_{1}k}*b_{kj_{1}}$ and $a_{i_{2}k}*b_{kj_{2}}$ present in $H_{i}$, will appear in the computation of $c_{i_{1}j_{1}}$ and $c_{i_{2}j_{2}}$. Thus, the processor cannot mix them in any way (since we are working with semirings). Therefore, the number of terms in a group should not exceed the memory of the processor. From this, we can conclude that
$t_{i}\leq O(r),\ \forall\ i$
and from above, we know that
$\sum\sqrt{t_{i}}\leq O(r)$
Thus, $\sum t_{i}=t$ is maximized when $t_{i}=\Theta(r)$ and the number of groups = $\Theta(\sqrt{r})$. Therefore, $t=O(n\sqrt{n})$ is the maximum number of terms that a processor can calculate in one round.
Figure 1: $n^{3}$ terms of matrix $C$ represented in the form of $n$ layers (each layer of distinct colour). One layer is obtained by multiplying one column of $A$ with one row of $B$. Each $H_{i}$ consists of terms only from one of the layers (here, three red filled cubes from Layer 1).
Theorem 2.5.
Matrix Multiplication in the MPC model with $n$ processors, each with $O(n)$ memory, requires $\Omega(\sqrt{n})$ rounds in semirings.
Proof 2.6.
As there are $n$ processors with $O(n)$ memory, a maximum of $\Theta(n^{2}\sqrt{n})$ terms can be calculated in one round. There are a total of $n^{3}$ terms in $C$. Hence, we need at least $\Omega(\sqrt{n})$ rounds to calculate the product matrix $C$.
From Theorem 2.1 and Theorem 2.5, we can conclude the following.
Theorem 2.7.
Square matrix multiplication in the MPC model with $n$ processors, each with $O(n)$ memory, requires $\Theta(\sqrt{n})$ rounds in semirings.
Let’s analyse the generalised case now. We have $n^{\alpha}$ processors where $0\leq\alpha\leq 2$. To keep the MPC model tight, each processor has $O(n^{2-\alpha})$ memory, so that the total memory is $O(n^{2})$.
Theorem 2.8.
Matrix multiplication in MPC model with $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, takes $\Theta(n^{\alpha/2})$ rounds for semirings.
Proof 2.9.
Similar to Theorem 2.1, we divide the matrices $A$, $B$, and $C$ into $n^{\alpha}$ sub-matrices of size $n^{1-\frac{\alpha}{2}}\times n^{1-\frac{\alpha}{2}}$. Let $M^{ij}$ and $p_{ij}$ be defined and distributed in the same way, except $i,j\in[n^{\frac{\alpha}{2}}]$ (instead of $\sqrt{n}$). We run the following algorithm in order to compute the matrix $C$.
for each $x$ from $1$ to $n^{\alpha/2}$ do
       Send elements of $A^{ix}$ and $B^{xj}$ to $p_{ij}$ $\forall\ i,\ j$
$p_{ij}$ locally computes $C^{ij}$ = $C^{ij}$ + $A^{ix}*B^{xj}$ $\forall\ i,\ j$
(* denotes Matrix Multiplication)
(Note that this step can be done without exceeding $O(n^{2-\alpha})$ memory)
Algorithm 2 Square Matrix Multiplication in $O(n^{\alpha/2})$ rounds
The above algorithm follows all the constraints of the MPC model. At the end of the algorithm, $p_{ij}$ contains $C^{ij}$ for all $1\leq i,j\leq n^{\alpha/2}$, meaning the entire product matrix $C$ has been calculated. Observing the argument presented in the proof of Theorem 2.5, in this setting, we can say that a processor can calculate at most $O(n^{3-3\alpha/2})$ terms in one round. Thus, a maximum of $O(n^{3-\alpha/2})$ terms can be calculated in total in one round. Hence, we need at least $\Omega(n^{\alpha/2})$ rounds to calculate the product matrix $C$. So, the given algorithm is the best possible algorithm for square matrix multiplication of general semirings in the MPC model.
@@ 3 Rectangular Matrix Multiplication
In this section, we consider the case when matrices $A$ and $B$ are rectangular matrices. There can be many types of instances in rectangular matrix multiplication. In this paper, we focus on two primary cases of type $(n,d,n)$ and $(d,n,d)$ and analyse them separately.
3.1 (n,d,n) Matrix Multiplicaton
In this case, we assume that we have $n$ processors with $O(n)$ memory each. Notice that in this case the matrices $A$ and $B$ can be stored in $O(nd)$ memory; however, the resulting $C$ matrix is a square matrix of size $n\times n$. Thus, we limit the total memory of the system to $O(n^{2})$. We divide the matrix $C$ into $n$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$ each, denoted by $C^{ij}$ (similar to Section 2) where $i,j\in[\sqrt{n}]$. Assign the computation of $C^{ij}$ to processor $p^{ij}$. Without loss of generality, assume that we scale the value of $d$ such that $d:=(\lceil\frac{d}{\sqrt{n}}\rceil\sqrt{n}).$ Divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{n}\times\sqrt{n}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,\ j\in[\sqrt{n}]$ and $q\in[\frac{d}{\sqrt{n}}]$. Observe that
We can assume that each processor $p_{ij}$ contains one row and one column of $A$ and $B$ respectively . Note that an element of $A$ needs to be sent to at most $\sqrt{n}$ processors. Thus, $p_{ij}$ needs to send $d\sqrt{n}$ elements. In one round, it can send $n$ elements and thus, in $\frac{d}{\sqrt{n}}$ rounds, it can send all the elements to the destination processors. Now let us analyse how many elements of $A$ and $B$ are received by each processor. $A^{iq}$ and $B^{qj}$ are $\sqrt{n}\times\sqrt{n}$ matrices containing a total of $n$ elements. Therefore, $p_{ij}$ needs a total of $(\frac{d}{\sqrt{n}}\cdot n)=d\sqrt{n}$ elements of $A$ and $B$ in order to compute $C^{ij}$. In one round, $p_{ij}$ can receive at most $O(n)$ elements and thus, in $\frac{d\sqrt{n}}{n}=\frac{d}{\sqrt{n}}$ rounds, it can receive all the elements that it requires for the computation of $C^{ij}$. Hence, we get the following simple algorithm.
for each $q$ from $1$ to $\frac{d}{\sqrt{n}}$ do
       receive elements of $A^{iq}$ and $B^{qj}$ from processors containing them
$p_{ij}$ locally computes $C^{ij}$ = $A^{iq}*B^{qj}$
Algorithm 3 $(n,d,n)$ Matrix Multiplication
The lower bound follows from Lemma 2.3. There are $n^{2}d$ terms that need to be calculated. In one round, a total of $n^{2}\sqrt{n}$ terms can be calculated. Therefore, $\Omega(\frac{d}{\sqrt{n}})$ rounds will be required. We proved the following theorem.
Theorem 3.1.
Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ respectively, with $n$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings, where $d\leq n$.
@@
In the $(d,n,d)$ case, the size of all three matrices is bounded by $nd$; thus, it is sufficient to have a total memory of size $O(nd)$. We further divide this case into two subcases: (i) when there are $n$ processors each with $O(d)$ memory, and (ii) when there are $d$ processors each with $O(n)$. First, we prove the following two lemmas that will help us to analyse both these cases.
@@
Lemma 3.2.
If there are $t$ processors each containing matrices of size $\sqrt{k}\times\sqrt{k}$, and the memory of each processor is $O(k)$, then the sum of these matrices can be computed in $O(\log_{k}t)$ rounds in the MPC model.
Proof 3.3.
Suppose $M_{1},M_{2},\ldots,M_{t}$ are the $t$ matrices and $X=\sum M_{l}$. Let $m_{ij}^{l}$ denote the element at $i^{th}$ row and $j^{th}$ column of $M_{l}$. Therefore $x_{ij}=\sum m_{ij}^{l}$. We divide the $t$ processors into $k$ groups of $(\frac{t}{k})$ processors such that each group is responsible for computing one $x_{ij}$. Let’s just focus on the first group, the same algorithm will be followed in the other groups. Let $p_{1},p_{2},\ldots,p_{t/k}$ be the processors in the first group. In the first round, all $t$ processor exchange messages such that $p_{1}$ receives elements $m_{11}^{1},m_{11}^{2},\ldots m_{11}^{k}$; $p_{2}$ receives elements $m_{11}^{k+1},m_{11}^{k+2},\ldots m_{11}^{2k}$ and so on (see Figure 2 ). Now, each $p_{i}$ computes the sum of all the elements it receives. Notice that all the values that are required to compute $x_{11}$ are present in $t/k$ processors, such that each processor contains only one value (the sum it has calculated). Now, in the next round, processor $p_{2},p_{3},\ldots p_{k}$ will send all the intermediate values they have calculated in the previous round to $p_{1}$. Similarly, $p_{k+2},p_{k+3},\ldots p_{2k}$ will send all the intermediate values they have calculated in the previous round to $p_{k+1}$ and so on (see Figure 3for reference). Now, notice that all the values that are required to compute $x_{11}$ are now present in the $(\frac{n}{k^{2}})$ processors, and each contains only one value. Repeat the previous step for $O(\log_{k}t)$ rounds. After $O(\log_{k}t)$, the final value of $x_{11}$ will be present in one processor. The same thing will happen in the other groups. Therefore, after $O(\log_{k}t)$ rounds processors will calculate $X$. This finishes the proof of the lemma. Notice that in the case where $t\leq k$, the algorithm will terminate in merely a single round. Therefore, in the above proof, we can assume $t>k$.
@@
Lemma 3.4.
If there are $t$ processors each with $O(k)$ memory such that each processor contains one element, then $\Omega(\log_{k}t)$ rounds are required to compute the sum of these elements in the MPC model, for semirings.
Proof 3.5.
To compute the sum, the only thing we can do is add the elements locally during a round of internal computation, and then send or receive this local sum(s) in some fashion during a communication round. We try to follow $x$–the number of processors that hold elements yet to be added together to obtain the final sum. Initially, $x$ is equal to $t$, and we need to reduce it to one in order to add all the elements. In one communication round, this $x$ can never decrease by more than a factor of $O(k)$, as one processor can only receive information from $O(k)$ processors during a communication round, due to its memory constraint. Thus, to reduce $x$ to one (which is the same as adding all elements), we will require $\Omega(\log_{k}t)$ communication rounds.
@@
3.2.1 n Processors with O(d) Memory
Let us divide the matrix $C$ into $d$ sub-matrices of size $\sqrt{d}\times\sqrt{d}$ each, denoted by $C^{ij}$, where $i,j\in[\sqrt{d}]$. We assign $(\frac{n}{d})$ processors to evaluate each sub-matrix. Also divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{d}\times\sqrt{d}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,j\in[\sqrt{d}]$ and $q\in[\frac{n}{\sqrt{d}}]$. We know that,

### v2

@@ Abstract
In this paper, we present algorithms to solve matrix multiplication problems in the MPC model. In particular, we consider the problem under various processor/memory constraints in the MPC model and prove the following results.
@@
    Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ ( where $d\leq n$) respectively can be done in,
    1. (a)
      $O(\sqrt{d}+\log_{d}n)$ rounds with $n$ processors and $\Theta(d)$ memory per processor
    2. (b)
      $O(\frac{d}{\sqrt{n}})$ rounds with $d$ processors and $\Theta(n)$ memory per processor.
@@
    Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ (where $d\leq n$) respectively, with $n$ processors of $\Theta(n)$ memory per processor can be done in $O(\frac{d}{\sqrt{n}})$ rounds.
@@
    The multiplication of two $d$-sparse matrices (matrices that contain at most $d$-nonzero elements in each row and in each column) with $n$ processors and $\Theta(d)$ memory per processor can be done in $O(d^{0.9})$ rounds.
@@ 1 Introduction
We are given two matrices $A$ and $B$, and our task is to compute $C=A\times B$. Matrix multiplication has been extensively studied in the centralized setting, leading to a long line of work on improving the asymptotic complexity of sequential algorithms (see, e.g., [12, 5, 6, 13, 11, 1]). However, these algorithms are not efficiently parallelizable due to the inherent dense communication pattern in them. This limitation has motivated the investigation of matrix multiplication in distributed and parallel computational models, including the clique model, low-bandwidth models, and related frameworks [10, 4, 3, 2, 7, 8]. Building on this line of research, in this paper we explore matrix multiplication in the Massively Parallel Computation (MPC) model, a framework that captures large-scale parallel computation with restricted memory per machine. We consider the problem $C=A\times B$ in the following three settings: when matrices $A$ and $B$ are (i) rectangular matrices of dimension $d\times n$ and $n\times d$, respectively, (ii) rectangular matrices of dimension $n\times d$ and $d\times n$, respectively and, (iii) $d$- sparse matrices i.e. both are $n\times n$ size matrices such that each matrix contains at most $d$ non-zero elements in each row and each column. For completeness, we provide a brief overview of the MPC model in section˜1.1, followed by a precise description of the computational settings in which we address these three variants of the matrix multiplication problem.
@@ 1.1 MPC Model
In this paper, we work with the massively parallel computation model (MPC) [9]. Given any problem of size (input ++ output) $m$ words, such that $m$ is too big to store on a single computer, in the MPC model, we assume that the problem is distributed over a fully connected network of computers (or processors). The computation proceeds in synchronous rounds. Processors perform computation locally, and at the end of a round, they exchange their messages. However, unlike some other models of distributed and parallel computation, in the MPC model, we assume that each computer has a limited memory. In particular, we assume that each computer has $\Theta(m^{\delta})$ memory and the number of computers in the network is $\Theta(m^{1-\delta})$, where $0<\delta<1$. Therefore, the total memory of the system is $\Theta(m)$ sufficient to store the given problem. Since each computer has $\Theta(m^{\delta})$ memory, it is allowed to send and receive $O(m^{\delta})$ words in each round. If any processor receives more than $\Theta(m^{\delta})$ words during the course of an algorithm, the algorithm is considered failed.
@@ 1.2 Setup
We assume that initially, matrices $A$ and $B$ are distributed evenly among all the processors. The structure of the distribution does not matter because we can reach any desired even distribution in one round (even if we start with an uneven distribution). At the end of the algorithm the elements of the output matrix $C$ should also be evenly distributed. We analyse the problem in the following cases.
  1. (i)
    $(d,n,d)$: when $A$ and $B$ are rectangular matrices size $d\times n$ and $n\times d$ respectively, where $d\leq n$.
  2. (ii)
@@
  3. (iii)
    $d$-sparse: AA and $B$ are $d$-sparse $n\times n$ matrices, i.e. each row and column of matrices $A$ and $B$ contains at most $d$ non-zero elements. Also we are interested in at most $d$ nonzero entries in each row and each column in $C$
@@ 1.3 Our Results
$(d,n,d)$:
In this case, both the input and output matrices have size $O(dn)$. Accordingly, it suffices to assume that the total memory available in the system is $\Theta(dn)$. We analyze this instance under two different processor–memory configurations:
  1. (i)
    For $n$ processor each with $\Theta(d)$ memory we design an algorithm with round complexity $O(\sqrt{d}+\log_{d}n)$
  2. (ii)
    For $d$ processor each with $\Theta(n)$ memory we design an algorithm with round complexity $(\tfrac{d}{\sqrt{n}})$ rounds.
$(n,d,n)$:
Here, the input matrices have size $O(nd)$, while the output matrix may be as large as $O(n^{2})$, corresponding to an $n\times n$ matrix. Therefore, we analyse this case under the configuration of $n$ processors, each with $\Theta(n)$ memory, and present an algorithm with round complexity $O(\tfrac{d}{\sqrt{n}})$.
$d$-sparse case:
Finally, we consider the case where both $A$ and $B$ are $n\times n$ matrices, each with at most $d$ nonzero entries per row and per column. Since we are interested in $d$ elements in the output matrix, the size of both input and output matrices is bounded by $O(nd)$. Therefore, we analyse this case with $n$ processors each with $\Theta(d)$ memory. The state-of-the-art for this is a trivial algorithm that takes $O(d)$ rounds. We improve that by giving an $O(d^{0.9})$ round algorithm.
1.4 High-level Idea of Our Technique
@@
In rectangular cases, we observed that in order to reduce inter-processor communication, the best case is when all processors try to compute a square submatrix of $C$. Thus, we divide the matrices $A$ and $B$ into smaller square sub-matrices of equal size depending on the local memory of the processor, and assign each submatrix to one of the processors. At the end, each processor needs to produce the corresponding submatrix of $C$. For the $d$-sparse case, we adapt the idea of Gupta et al.[7]. We first refine the analysis given in their paper and improve upon their results. Here, we would like to highlight that we improve the bound present in their paper from $O(d^{1.927})$ to $O(d^{1.925})$ for solving $d$-sparse matrix multiplication in the low bandwidth model. We do not explicitly mention that result here because that result is about the low-bandwidth model, and in this paper, we stick to the MPC model. However, the improved bound can be directly derived from the lemma˜3 that we prove in this paper (by appropriately setting the value of $\epsilon$ for the low bandwidth model). Then we use lemma˜3 to prove our result for $d$-sparse matrix multiplication in the MPC model. In their paper, they show that $d$-sparse matrix multiplication is equivalent to processing triangles in a $d$-degree graph. In other words, processing triangles is equivalent to a product $a_{ij}b_{jK}$, where $a_{ij}$ and $b_{jK}$ are some elements of matrices $A$ and $B$, respectively. They show that given any $d$-degree graph, say $T$, we can divide the graph into two parts such that one part, say $T^{\prime}$, is clustered - it contains layers of subgraphs of the original graph such that in each layer we have disjoint clusters of triangles which are small and dense, and the other part, say $T^{\prime}$, contains a low number of triangles. Now, in order to process all the triangles, we can first process all the triangles in the clustered part layer-by-layer using square matrix multiplication, then compute all the triangles $T^{\prime}$ using brute force. We noticed that there was some suboptimality in bring down the unprocessed triangles from $T$ to $T^{\prime}$. In this paper, we remove that suboptimality and then use that result as a black box to solve $d$-sparse matrix multiplication in the MPC model.
@@ 2 Rectangular Matrix Multiplication
In this section, we consider the case when matrices $A$ and $B$ are rectangular matrices. There can be many types of instances in rectangular matrix multiplication. In this paper, we focus on two primary cases of type $(d,n,d)$ and $(n,d,n)$ and analyse them separately in subsequent sections.
@@
In the $(d,n,d)$ case, the size of all three matrices is bounded by $nd$; thus, it is sufficient to have a total memory of size $O(nd)$. We analyse this case under two different memory constraints, two subcases: (i) when there are $n$ processors each with $O(d)$ memory, and (ii) when there are $d$ processors each with $O(n)$, where $d\leq n$. First, we prove the following two lemmas that will help us to analyse both these cases.
@@
Lemma 1
If there are $t$ processors each containing matrices of size $\sqrt{k}\times\sqrt{k}$, and the memory of each processor is $\Theta(k)$, then the sum of these matrices can be computed in $O(\log_{k}t)$ rounds in the MPC model.
Proof
Suppose $M_{1},M_{2},\ldots,M_{t}$ are the $t$ matrices and $X=\sum M_{l}$. Let $m_{ij}^{l}$ denote the element at $i^{th}$ row and $j^{th}$ column of $M_{l}$. Therefore $x_{ij}=\sum m_{ij}^{l}$. We divide the $t$ processors into $k$ groups of $(\frac{t}{k})$ processors such that each group is responsible for computing one $x_{ij}$. Let’s just focus on the first group, the same algorithm will be followed in the other groups. Let $p_{1},p_{2},\ldots,p_{t/k}$ be the processors in the first group. In the first round, all $t$ processor exchange messages such that $p_{1}$ receives elements $m_{11}^{1},m_{11}^{2},\ldots m_{11}^{k}$; $p_{2}$ receives elements $m_{11}^{k+1},m_{11}^{k+2},\ldots m_{11}^{2k}$ and so on (see fig.˜1 ). Now, each $p_{i}$ computes the sum of all the elements it receives. Notice that all the values that are required to compute $x_{11}$ are present in $t/k$ processors, such that each processor contains only one value (the sum it has calculated). Now, in the next round, processor $p_{2},p_{3},\ldots p_{k}$ will send all the intermediate values they have calculated in the previous round to $p_{1}$. Similarly, $p_{k+2},p_{k+3},\ldots p_{2k}$ will send all the intermediate values they have calculated in the previous round to $p_{k+1}$ and so on (see fig.˜2 for reference). Now, notice that all the values that are required to compute $x_{11}$ are now present in the $(\frac{n}{k^{2}})$ processors, and each contains only one value. Repeat the previous step for $O(\log_{k}t)$ rounds. After $O(\log_{k}t)$, the final value of $x_{11}$ will be present in one processor. The same thing will happen in the other groups. Therefore, after $O(\log_{k}t)$ rounds processors will calculate $X$. This finishes the proof of the lemma. Notice that in the case where $t\leq k$, the algorithm will terminate in merely a single round. Therefore, in the above proof, we can assume $t>k$.
@@
2.1.1 n Processors with $O(d)$ Memory:
Since $C$ is of size $d\times d$, let us divide $C$ into $d$ sub-matrices of size $\sqrt{d}\times\sqrt{d}$ each, denoted by $C^{ij}$, where $i,j\in[\sqrt{d}]$. We assign $(\frac{n}{d})$ processors to evaluate each sub-matrix. Also divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{d}\times\sqrt{d}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,j\in[\sqrt{d}]$ and $q\in[\frac{n}{\sqrt{d}}]$. We know that,